# Coprime Bivariate Bicycle Codes and their Properties

### v1

@@ Abstract
This work (1) proposes a novel numerical algorithm to accelerate the search process for good Bivariate Bicycle (BB) codes and (2) defines a new variant of BB codes suitable for quantum error correction. The proposed acceleration search algorithm reduces the search space by excluding some equivalent codes from the search space, as well as setting thresholds to drop bad codes at an early stage. A number of new BB codes found by this algorithm are reported. The proposed variant of BB codes employs coprimes to construct groups via polynomials as the basis for the BB code, rather than using the standard BB codes with unconstrained constructors. In contrast to vanilla BB codes, where parameters remain unknown prior to code discovery, the rate of the proposed code can be determined beforehand by specifying a factor polynomial as an input to the numerical search algorithm. Using this coprime BB construction, we found a number of surprisingly short to medium-length codes that were previously unknown.
@@ 1 Introduction
This challenge motivated us to develop an algorithm that accelerates the search for good BB codes and even constructs codes with pre-determined parameters. In this work, we propose a fast numerical algorithm to search for good BB codes by excluding certain polynomial combinations. In addition, a new construction of BB codes is proposed that allows us to customize the code rate before performing a search, much in contrast to prior search techniques that identified the rate only after returning a new code as a search result. This new method involves selecting two coprime numbers and a factor polynomial, leading us to name this variant coprime-BB codes.
@@ 2.1 Stabilizer Codes and CSS Codes
$H_{CSS}=\begin{bmatrix}H_{X}&\bm{0}\\ \bm{0}&H_{Z}\end{bmatrix},$ (3)
@@ 3.1 Code equivalence
Next, we will prove that $C_{1}$ and $C_{2}$ have the same distance. Let $C_{l}$ be the anti-diagonal matrix of size $l\times l$. We have $C_{l}C_{l}=I$ and $C_{l}MC_{l}=M^{T}$ for any given matrix $M$ of the same size. As
$[A^{T}|B^{T}]=C_{lm}[A|B]\begin{bmatrix}C_{lm}&0\\ 0&C_{lm}\end{bmatrix}$ (11)
, we know $[A^{T}|B^{T}]$ and $[A|B]$ have the same rank and $k_{1}=k_{2}$. Let $Z^{\bm{p}}$ be an arbitrary logical $Z$ operator of $C_{2}$, where $\bm{p}=(\bm{\alpha}|\bm{\beta})$ is a length $n$ binary vector and $\bm{\alpha},\bm{\beta}$ are binary vectors of length $n/2$. By definition, we have $[A^{T}|B^{T}]\bm{p}^{T}=A\bm{\alpha}^{T}+B\bm{\beta}^{T}=\bm{0}$. I.e.,
$C_{lm}AC_{lm}\bm{\alpha}^{T}+C_{lm}BC_{lm}\bm{\beta}^{T}=\bm{0}$ (12)
. Recall that $C_{lm}C_{lm}=I$. By multiplying both sides with $C_{lm}$, we get
$AC_{lm}\bm{\alpha}^{T}+BC_{lm}\bm{\beta}^{T}=\bm{0}$ (13)
, and we know that $(\bm{\alpha}C_{lm}^{T}|\bm{\beta}C_{lm}^{T})$ is a logical $Z$ operator of $C_{1}$ with the same weight of $\bm{p}$. Therefore, $d_{1}\leq d_{2}$ and, similarly, $d_{2}\leq d_{1}$. Thus, $d_{2}=d_{1}$.∎
@@ 3.2 Code Selection
This condition is equivalent to decoding a code with an enlarged parity-check matrix $H^{\prime}=\begin{bmatrix}H_{Z}\\ \bm{u}\end{bmatrix}$ and syndrome $\bm{s}=\begin{bmatrix}\bm{0}\\ 1\end{bmatrix}$. In [4], the authors used the BP-OSD [9] algorithm to perform the decoding. The distance can be tightly bounded after multiple rounds of iterations. Since we are finding codes with high distances, this process can be aborted if the upper bound is lower than the desired distance $\tau_{d}$.
@@ 4 Coprime Bivariate Bicycle Codes
By adapting the new construction to Algorithm 1, we propose Algorithm 2. Using Algorithm 2, we found a number of interesting coprime-BB codes, which are shown in Table 2. The results are not exhaustive as numerous polynomials meet the condition $\text{GCD}(a(\pi),b(\pi))=p(\pi)$. As the algorithm outputs codes with the same $k$, the results shown were obtained by running the search algorithm for a short duration and selecting the output code with the highest $d$. Additionally, we visualized the newly found codes along with the codes proposed in [4] in Fig. 1 as a 3D plot of the $n,k,d$ dimensions. “Better codes” are in the red subspace, which have higher $k$ and $d$ relative to their lower $n$ values, i.e., their cost in physical qubits in balanced by the logical qubits and sufficiently large distance for error correction.
We observe that the $[[126,12,10]]$ code offers the highest $d$ and one of the highest $k$, which is comparable to the $[[144,12,12]]$ “gross” code proposed in [4], making it suitable for scenarios where error rates are moderate to high, and qubit resources are ample. The $[[42,6,6]]$ code has a low $d$ but provides the highest rate in these codes and is best suited for scenarios with low error rates or where hardware resources are limited. The $[[70,6,8]]$ code offers a balanced trade-off between error correction capability and code length, making it a versatile option for environments with moderate physical error rates and resource constraints.
@@ 6 Related Work
While conducting this work, various constructions based on BB codes have been proposed to achieve different objectives. Voss et al. [11] expanded the concept of BB codes by introducing an additional type of indeterminate variable, leading to the creation of trivariate bicycle (TB) codes. These new codes reduce the weight of stabilizers from 6 to 4-5, making them more practical for hardware implementation. However, it is important to note that some of these codes exhibit a lower rate or distance compared to BB codes of similar length.
@@ 7 Conclusion and Future Work
We developed algorithms for fast numerical searches for the discovery of BB codes. Furthermore, we proposed a novel construction of BB codes by choosing a factor polynomial from $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$, where $l$ and $m$ are coprime integers. The new construction enables us to know the rate of BB codes before constructing them. Simulations have been done to illustrate the error rates of the newly found codes. Our $[[126,12,10]]$ code achieves a slightly higher error rate than the $[[144,12,12]]$ code with fewer qubits. However, a challenge that remains is the practical implementation of these codes, specifically in mapping them onto quantum architectures that are constrained by the limitations of current quantum device technologies.

### v2

@@ Abstract
This work (1) proposes a novel numerical algorithm to accelerate the search process for good Bivariate Bicycle (BB) codes and (2) defines a new subclass of BB codes suitable for quantum error correction. The proposed acceleration search algorithm reduces the search space by excluding some equivalent codes from the search space, as well as setting thresholds to drop bad codes at an early stage. A number of new BB codes found by this algorithm are reported. The proposed subclass of BB codes employs coprimes to construct groups via polynomials as the basis for the BB code, rather than using the standard BB codes with unconstrained constructors. In contrast to vanilla BB codes, where parameters remain unknown prior to code discovery, the rate of the proposed code can be determined beforehand by specifying a factor polynomial as an input to the numerical search algorithm. Using this coprime BB construction, we found a number of surprisingly short to medium-length codes that were previously unknown.
@@ 1 Introduction
This challenge motivated us to develop an algorithm that accelerates the search for good BB codes and even constructs codes with pre-determined parameters. In this work, we propose a fast numerical algorithm to search for good BB codes by excluding certain polynomial combinations. In addition, a new construction of BB codes is proposed that allows us to customize the code rate before performing a search, much in contrast to prior search techniques that identified the rate only after returning a new code as a search result. This new method involves selecting two coprime numbers and a factor polynomial, leading us to name this subclass coprime-BB codes.
@@ 2.1 Stabilizer Codes and CSS Codes
$H_{CSS}=\begin{pmatrix}H_{X}&\bm{0}\\ \bm{0}&H_{Z}\end{pmatrix}\leavevmode\nobreak\ ,$ (3)
@@ 3.1 Code equivalence
Next, we will prove that $C_{1}$ and $C_{2}$ have the same distance. Let $C_{l}$ be the anti-diagonal matrix of size $l\times l$. We have $C_{l}C_{l}=I$ and $C_{l}MC_{l}=M^{T}$ for any given matrix $M$ of the same size. As per
$[A^{T}|B^{T}]=C_{lm}[A|B]{\begin{bmatrix}C_{lm}&0\\ 0&C_{lm}\end{bmatrix},}$ (11)
we know $[A^{T}|B^{T}]$ and $[A|B]$ have the same rank and $k_{1}=k_{2}$. Let $Z^{\bm{p}}$ be an arbitrary logical $Z$ operator of $C_{2}$, where $\bm{p}=(\bm{\alpha}|\bm{\beta})$ is a length $n$ binary vector and $\bm{\alpha},\bm{\beta}$ are binary vectors of length $n/2$. By definition, we have $[A^{T}|B^{T}]\bm{p}^{T}=A\bm{\alpha}^{T}+B\bm{\beta}^{T}=\bm{0}$. I.e.,
$C_{lm}AC_{lm}\bm{\alpha}^{T}+C_{lm}BC_{lm}\bm{\beta}^{T}=\bm{0}.$ (12)
Recall that $C_{lm}C_{lm}=I$. By multiplying both sides with $C_{lm}$, we get
$AC_{lm}\bm{\alpha}^{T}+BC_{lm}\bm{\beta}^{T}=\bm{0},$ (13)
and we know that $(\bm{\alpha}C_{lm}^{T}|\bm{\beta}C_{lm}^{T})$ is a logical $Z$ operator of $C_{1}$ with the same weight of $\bm{p}$. Therefore, $d_{1}\leq d_{2}$ and, similarly, $d_{2}\leq d_{1}$. Thus, $d_{2}=d_{1}$.∎
@@ 3.2 Code Selection
This condition is equivalent to decoding a code with an enlarged parity-check matrix $H^{\prime}=\begin{bmatrix}H_{Z}\\ \bm{u}\end{bmatrix}$ and syndrome $\bm{s}=\begin{bmatrix}\bm{0}\\ 1\end{bmatrix}$. In [4], the authors used the BP-OSD [9, 10] algorithm to perform the decoding. The distance can be tightly bounded after multiple rounds of iterations. Since we are finding codes with high distances, this process can be aborted if the upper bound is lower than the desired distance $\tau_{d}$.
@@ 4 Coprime Bivariate Bicycle Codes
Coprime-BB codes can be viewed as a special case of generalized bicycle (GB) codes. In GB codes, polynomials are identified as sums of cyclic shift matrices. In contrast, coprime-BB codes identify polynomials as sums of the Kronecker products of two cyclic shift matrices with coprime dimensions. By adapting the new construction to Algorithm 1, we propose Algorithm 2. The latter algorithm results in a significantly reduced search space as only coprimes and codes with desired $k$ are being considered. Using Algorithm 2, we found a number of interesting coprime-BB codes, which are shown in Table 2. The results are not exhaustive as numerous polynomials meet the condition $\text{GCD}(a(\pi),b(\pi))=p(\pi)$. As the algorithm outputs codes with the same $k$, the results shown were obtained by running the search algorithm for a short duration and selecting the output code with the highest $d$. Additionally, we visualized the newly found codes along with the codes proposed in [4] in Fig. 1 as a 3D plot of the $n,k,d$ dimensions. “Better codes” are in the red subspace, which have higher $k$ and $d$ relative to their lower $n$ values, i.e., their cost in physical qubits in balanced by the logical qubits and sufficiently large distance for error correction.
We observe that the $[[126,12,10]]$ code offers the highest $d$ and one of the highest $k$, which is comparable to the $[[144,12,12]]$ “gross” code proposed in [4], making it suitable for scenarios where error rates are moderate to high, and qubit resources are ample. We also noted that in [10], the author found a [[$126,12,10$]] with different polynomials, i.e., we are not the first to discover this particular code, yet our search algorithm identified it based on a different basis. The $[[42,6,6]]$ code has a low $d$ but provides the highest rate in these codes and is best suited for scenarios with low error rates or where hardware resources are limited. The $[[70,6,8]]$ code offers a balanced trade-off between error correction capability and code length, making it a versatile option for environments with moderate physical error rates and resource constraints.
@@ 6 Related Work
Panteleev and Kalachev proposed a BP-OSD decoder and discovered the [[$126,12,10$]] code [10]. More recently, various constructions based on BB codes have been proposed to achieve different objectives. Voss et al. [11] expanded the concept of BB codes by introducing an additional type of indeterminate variable, leading to the creation of trivariate bicycle (TB) codes. These new codes reduce the weight of stabilizers from 6 to 4-5, making them more practical for hardware implementation. However, it is important to note that some of these codes exhibit a lower rate or distance compared to BB codes of similar length.
@@ 7 Conclusion and Future Work
We developed algorithms for fast numerical searches for the discovery of BB codes. Furthermore, we proposed a novel construction of BB codes by choosing a factor polynomial from $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$, where $l$ and $m$ are coprime integers. The new construction enables us to know the rate of BB codes before constructing them. Simulations have been done to illustrate the error rates of the newly found codes. The $[[126,12,10]]$ code achieves a slightly higher error rate than the $[[144,12,12]]$ code with fewer qubits. However, a challenge that remains is the practical implementation of these codes, specifically in mapping them onto quantum architectures that are constrained by the limitations of current quantum device technologies.